There are four categories. We like to test whether P(A) = 0.4, P(B) =0.3, P(C) = 0.15, and P(D) = 0.15 with

0.05 alpha.

Category A B C   D

Frequency 250 100 70 60

If the test statistics = 33.02, which one is correct? Choose all applied.

a.

Upper cv = 7.814

b.

Upper cv = 9.488

c.

p-value = 1.183271e-06

d.

p-value is 3.189539e-07
Answer #1

The solution to this problem is given by

Sretion Number of Categories = nay .. Degence of freedom= ran-1=4 -1 =3 level of siguificance = x=0.05 Test Statistica x2 33.